Circular No. 8674

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ae, 2006af, 2006ag, 2006ah, 2006ai Several apparent supernovae have been discovered on CCD images

(all unfiltered except for the spectroscopic SDSS discoveries): 2006ae and 2006af by M. SubbaRao and the Sloan Digital Sky Survey (SDSS) collaboration (cf. IAUC 8668, CBET 400; magnitudes below are estimated r-band values; both objects at center of their host galaxies); 2006ag and 2006ah on Palomar NEAT survey images via the “Nearby Supernova Factory” collaboration (details on CBET 402); and 2006ai by P. Luckas, O. Trondal, and M. Schwartz (cf. IAUC 8668). SNe 2006ae, 2006af, and 2006ah all appear to be type-Ia supernovae within a weak of maximum light upon discovery, while 2006ag is a type-IIn event. Additional approximate magnitudes of 2006ai in ESO 5-G9: 2005 Dec. 16.79 UT, [18.5; 2006 Feb. 19.52, 16.0.

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ae Feb. 2.51 14^{h}48^{m}23.^{s}27 +21^{o}47′51′′.5 20.7  —
2006af Feb. 4.38 10 33 57.18 +20 20 25.7 19.5  —
2006ag Feb. 9.4 12 00 17.12 +28 36 22.0 17.9  —
2006ah Feb. 9.6 13 46 13.71 − 9 07 50.6 18.6  —
2006ai Feb. 17.54 7 29 52.16 −84 02 20.5 16.2 2′′.5 W, 1′′.8 S

GRB 060218 = SUPERNOVA 2006aj A. Soderberg, California Institute of Technology; E. Berger, Observato-

ries of the Carnegie Institution of Washington; and B. Schmidt, Australian National University, report that a spectrogram (range 320–1000 nm) of GRB 060218 (cf. Cusumano et al., GCN Observation Report Circ. 4775; α = 3^{h}21^{m}37^{s}, δ = +16^{o}51′58′′, equinox 2000.0), obtained with Gemini-South telescope (+ GMOS) on Feb. 21.024 UT, shows that underlying a power-law continuum are features consistent with a broad-lined type-Ib/c supernova (designated 2006aj) near maximum light, confirming the findings of Masetti et al. (GCN 4803). The Gemini-South spectrum also confirms, from several narrow emission lines, the redshift of z = 0.033 (Mirabal, who reported r = 17.6 on Feb. 19.1438 via GCN 4784, and who notes a preoutburst extended object at the GRB location via GCN 4783). Marshall et al. provide a precise position of the optical counterpart of GRB 060218 from images obtained with the Swift UltraViolet and Optical Telescope: α = 3^{h}21^{m}39^{s}.71, δ = +16^{o}52′02′′.6 (estimated 1σ error ∼ 1′′.0; cf. GCN 4779). Zheng et al. report (on GCN 4802) the following R magnitudes for the optical counterpart of GRB 060218: Feb. 18.4522, 17.8; 20.4471, 18.2.

2006 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green